OpenStudy (anonymous):

what is the solution set for the equation x^2-3x=10

OpenStudy (mathstudent55):

Subtract 10 to the left side. Then try factoring the left side.

OpenStudy (anonymous):

x^2 - 3x=10 x^2-3x-10=0 (x-5)(x+2)=0

OpenStudy (texaschic101):

x^2 - 3x = 10 x^2 - 3x - 10 = 0 (x - 5)(x + 2) = 0 x - 5 = 0 x = 5 + 0 x = 5 and x + 2 = 0 x = -2 + 0 x = -2 answer : x = 5 and x = -2

OpenStudy (anonymous):

Note that this is called the zero product rule, where we equate 0 to something like (x-5)(x+2), and thus from here we can conclude that both sides are equal to zero. We can't do this if for example (x-3)(x-4)=5 though :) have a nice day!